# Toward Generalist Semi-supervised Regression via Decoupled Representation Distillation

### v1

@@ Abstract
Semi-supervised regression (SSR), which aims to predict continuous scores of samples while reducing reliance on a large amount of labeled data, has recently received considerable attention across various applications, including computer vision, natural language processing, and audio and medical analysis. Existing semi-supervised methods typically apply consistency regularization on the general regression task by generating pseudo-labels. However, these methods heavily rely on the quality of pseudo-labels, and direct regression fails to learn the label distribution and can easily lead to overfitting. To address these challenges, we introduce an end-to-end Decoupled Representation distillation framework (DRILL) which is specially designed for the semi-supervised regression task where we transform the general regression task into a Discrete Distribution Estimation (DDE) task over multiple buckets to better capture the underlying label distribution and mitigate the risk of overfitting associated with direct regression. Then we employ the Decoupled Distribution Alignment (DDA) to align the target bucket and non-target bucket between teacher and student on the distribution of buckets, encouraging the student to learn more robust and generalized knowledge from the teacher. Extensive experiments conducted on datasets from diverse domains—including audio, text, image, and medical data—demonstrate that the proposed DRILL has strong generalization and outperforms the state-of-the-art competing methods.
Introduction
@@
Semi-supervised regression (SSR) aims to rank or predict sample values using deep learning methods while reducing reliance on a large amount of labeled data. It has been widely applied in various scenarios, including age estimation, clinical score prediction, and audio-based assessments. With the rapid advancement of deep learning in recent years, regression tasks using deep learning models across various domains have attracted significant attention (Yang et al. 2024; Van Engelen and Hoos 2020). Some approaches directly employ models such as CNNs for images and RNNs for time series, optimizing them with loss functions like MSE or MAE on the available labeled data (Mohammadi Foumani et al. 2024; Yang et al. 2022), see Fig. 1 (a). However, this straightforward application often risks overfitting, see Fig. 1 (b), especially in scenarios with limited labeled samples, as the performance of deep learning models heavily relies on the availability of labeled data. In real-world applications, obtaining a large amount of labeled data across diverse domains remains a persistent challenge. (Gui et al. 2024)
Figure 1: (a) The framework of existing SSR methods, which typically directly predict the score using MAE is referred to as Direct Regression (DR) methods. (c) The framework DRILL establishes an end-to-end decoupled representation distillation framework. (b) and (d) are the training loss (MAE) and test SRCC curves for DR and DRILL on the UTKFace dataset (Zhang, Song, and Qi 2017). DR achieves a very low training loss, but its lower SRCC indicates the risk of overfitting. In contrast, DRILL maintains a high SRCC as training converges, demonstrating its better generalization.
@@
Therefore, several methods that assign the pseudo-labels on the unlabeled samples have been proposed to address this issue (Zhang et al. 2025; Zheng et al. 2025; Min, Bai, and Li 2024). Pseudo-labels are typically generated by training on the available labeled samples, and consistency regularization is often used to assist the optimization process. While pseudo-labeling helps mitigate the challenge of limited labeled data, overfitting may also occur, as incorrect labels are treated as true labels during training, the model is likely to overfit to the wrong annotations. The effectiveness of these approaches heavily relies on the quality of the pseudo-labels, and assessing the reliability of the generated pseudo-labels remains a challenging task.
To address these challenges, we introduce an end-to-end decouple distillation framework (DRILL), which is specially designed for SSR, see Fig. 1 (c). DRILL is implemented through Discrete Distribution Estimation (DDE), which transforms the regression task into a distribution estimation problem. In this approach, the continuous ground truth values are divided into several buckets, and the model predicts a probability distribution over these buckets. By dividing continuous labels into discrete levels, the model can capture the inherent structure of the labels, enabling it to model the ordinal and interval relationships among them. This approach also effectively mitigates the risk of overfitting, as it offers a smoother learning objective and richer label distribution than direct numerical regression, especially when the training data is limited (see Fig. 1 (d)).
To further enhance generalization, we introduce a novel decoupling distribution alignment (DDA) approach, where we decouple the distribution into two parts for alignment, rather than aligning the distribution directly. The discrete distribution is the probability of the score being located in this bucket, and the final score can be estimated by the expectation, making the student learn more robust and more generalizable knowledge representation provided by the teacher. Although there are some attempts by applying the decouple distillation on the classification task (Zhao et al. 2022; Li, Wang, and Cui 2023; Feng et al. 2024). There is no work on semi-supervised regression. And the main difference here is that we aim to leverage the decoupled distribution alignment to highlight the target buckets and non-target buckets during the distillation, and the targets are recognized by the ground truth and generated pseudo labels, which help extract more fine-grained features, thereby avoiding overfitting and instability that may arise from directly fitting noisy or complex ground truth. Through distillation, we leverage the distribution over buckets for learning, avoiding sole reliance on pseudo labels and reducing the negative impact of incorrect pseudo labels on the model. In summary, this work makes the following main contributions.
  • We for the first time built an end-to-end decouple distillation framework, DRILL, for semi-supervised regression, where we transform the direct regression into discrete distribution estimation (DDE) to mitigate the overfitting issue commonly encountered in direct regression.
@@
  • We further introduce a novel alignment method, decouple distribution alignment (DDA), which enables the student model to identify important buckets and mitigate the negative impact of noise. This facilitates the learning of richer and more accurate knowledge, thereby enhancing the generalization capability of the SSR task.
@@
  • Extensive experiments were conducted on the datasets from various domains, including audio, text, image, and medical data. The experimental results demonstrate that DRILL outperforms the state-of-the-art semi-supervised regression methods.
Related Works
Semi-supervised Learning
Semi-supervised Classification. Semi-supervised classification aims to improve classification performance by leveraging a small amount of labeled data along with a large amount of unlabeled data by generating the corresponding pseudo labels. The pseudo-label-based methods typically depend on the smoothness assumption that neighboring data points in the feature space tend to exhibit similar labels (Yang et al. 2022; Mey and Loog 2023; Berthelot et al. 2019; Sohn et al. 2020). The pesudo-label generation approach can effectively enhance the dataset and help the model learn fine-grained boundaries (Lee et al. 2013). As an effective way to refine the knowledge from existing methods, knowledge distillation is also used for pseudo generation, for example, Xie et al. present a pseudo-labeling approach based on knowledge distillation, using diverse student training techniques. This involves initially training a teacher model on labeled images to generate pseudo-labels for unlabeled examples (Xie et al. 2020b). Although some knowledge distillation methods have been applied to semi-supervised classification, their full potential has not been explored in the SSR task.
@@
Semi-supervised Regression SSR refers to predicting continuous real-valued outputs rather than discrete class labels using a dataset that contains only a small subset of labeled instances. The semi-supervised regression methods can be roughly categorized into pseudo-label-based methods, consistency regularization, and surrogate task-based methods. Jo et al. (Jo, Kahng, and Kim 2024) propose an SSR framework that uses uncertainty estimation to guide pseudo-labeling, alongside a pseudo-label calibration approach that enhances pseudo-label quality by propagating information from labeled to unlabeled samples. Semi-supervised deep kernel learning (Jean, Xie, and Ermon 2018) minimizes predictive variance on unlabeled data through consistency regularization, leveraging the robust uncertainty quantification provided by Gaussian processes. Some surrogate task-based methods exploit the ranking information among samples as auxiliary for regression (Qiao, Chen, and Zhu 2022; Huang, Fu, and Tsao 2024). For example, Rankup (Huang, Fu, and Tsao 2024) constructs pairwise ranking labels from continuous targets and applies standard semi-supervised classification techniques with an auxiliary ranking classifier. Although these methods have achieved remarkable success in SSR, most of them directly predict a single continuous value, which makes them sensitive to noise and prone to overfitting. Therefore, instead of operating directly on the regression output, we transform the SSR to a discrete distribution estimation to mitigate the risk of overfitting, thereby leading to better performance.
Distribution Alignment in Distillation
@@
Distribution alignment is a strategy used during student model training to match not only the outputs of the teacher model but also its underlying distributions. Existing approaches can be roughly categorized into logit-level alignment and feature-level alignment. Logit-level alignment aims to align the output distribution (e.g., softmax probabilities) of the teacher and student (Hinton, Vinyals, and Dean 2015; Zhao et al. 2022). For example, the multi-level logit distillation method (Jin, Wang, and Lin 2023) employs multi-level prediction alignment to facilitate instance prediction learning in the student model. Curriculum temperature (Li et al. 2023) introduces temperature scheduling for dynamically adjusting softening intensity across training phases, which smooths gradient flow. The feature-level distribution alignment matches the representation, which is often done by maximum mean discrepancy, contrastive, and correlation-based loss (Gou et al. 2021; Chen et al. 2021). Notable advancements in feature-level distillation encompass DiffKD (Huang et al. 2023), training diffusion denoisers with teacher features to enforce attention map congruence in critical regions; WCoRD (Chen et al. 2021), constructing global-local hybrid contrastive objectives with Wasserstein dual loss to align deep geometric structures. These alignment approaches have demonstrated effectiveness in knowledge transfer. However, they struggle to capture fine-grained information, resulting in limited generalization. To address this, instead of aligning representations directly as in prior work, we propose DDA, which leverages the decouple representation distillation to align the distributions over buckets, enabling the student to learn richer and more generalized knowledge.
Figure 2: The Overview of DRILL. (1) The input includes both labeled and unlabeled data, and the weak and strong augmentations are applied to the teacher and student models, respectively. The teacher model is exclusively trained on the labeled data and is then used to generate pseudo labels for the unlabeled samples. The student model is trained on the entire dataset, utilizing the ground-truth labels for labeled samples and the pseudo labels for the unlabeled ones. (2) The DDE transforms score prediction into a classification task over discrete buckets, effectively mitigating the risk of overfitting. (3) The DDA first identifies the target bucket to construct a binary probability distribution, then aligns the predicted distributions over buckets between the teacher and student models separately for the target and non-target parts. This decoupled alignment enables the student model to focus on critical information and learn more generalized representations.
Methodology
@@
Notations. We assume the data set $\mathcal{X}$ consists of $n$ labeled samples $\mathcal{X}_{l}=\{{({x}_{i},y_{i})}_{i=1}^{n}\}$ and $m$ unlabeled samples $\mathcal{X}_{u}=\{{({x}_{i})}_{i=m}^{n+m}\}$, with instance $x\in\mathbb{R}$ for image, text audio or tabular medical data. $\mathcal{Y}=\{y_{i}\}$ denotes the set of ground truth labels, where each label $y_{i}$ is a continuous value in $\mathbb{R}$.
@@
Problem Statement. Deep semi-supervised regression can be reformulated as given the training set $\mathcal{X}_{T}$ and $\mathcal{X}_{T}$ = $\mathcal{X}_{l}$, we aim to train a mapping function $f:\mathcal{X}_{u}\mapsto\mathcal{Y}$ which maps each input ${x}_{i}\in\mathcal{X}_{u}$ to a continuous scalar $y_{i}\in\mathbb{R}$. Some self-supervised methods extract features by capturing relationships between samples, and thus often include unlabeled data as input. Therefore, in this case, the $\mathcal{X}_{T}$ consists of $\mathcal{X}_{l}$ and $\mathcal{X}_{u}$.
The Overview of DRILL
@@
As shown in Fig. 2, the proposed DRILL predicts scores by establishing a decoupled representation distillation framework, where the ground truth is transformed into a distribution over several buckets by DDE, effectively converting the regression task into a discrete distribution estimation problem for both teacher and student models. We then apply DDA over buckets to enable the student model to learn fine-grained and more generalized knowledge.
@@
This distillation process is particularly effective for SSR task due to two key factors: (1) we reformulate the direct regression task as a discrete distribution estimation problem, which effectively mitigates the risk of overfitting of direct regression especially in the scenario that labeled samples are limited; and (2) we introduce a novel decoupled distribution alignment strategy for representation distillation over the buckets, enabling the model to capture fine-grained features more effectively than conventional alignment methods. Finally, the model is jointly trained by combining the loss functions of the student and teacher, augmented with the proposed decoupled distribution alignment.
Discrete Distribution Estimation (DDE)
@@
Conventional SSR methods typically predict the score by directly mapping the input sample ${x}_{i}$ to the target value $y$. However, this approach is highly sensitive to noise and irrelevant features, often resulting in poor generalization and a high risk of overfitting, especially when few labels are available. Therefore, inspired by (Gao et al. 2017; Xu, Liu, and Geng 2019), we define several buckets $\mathcal{B}=\{{\mathbf{b}_{i}}\}_{i=1}^{L}$ based on the range of ground truth $y$, converting the regression task into a discrete distribution estimation problem. The obtaining of $\mathcal{B}$ and the size $L$ depend on the $y$, if we assume the bucket capacity is $c$, $\mathbf{b}_{i}=(i-1)\times c$, where each score is assigned to one corresponding bucket. Then, $y$ can be obtained by calculating the expectation as follows:
@@
where $\tilde{p}_{i}$ is the output probability of the model with the corresponding bucket $\mathbf{b}_{i}$}. Therefore, the mapping function $f$ in SSR is modified to learn a distribution $\mathcal{P}=\{p_{i}\}_{i=1}^{L}$ over predefined buckets. By transforming the numerical regression into an interval prediction problem, DDE offers a smoother learning objective and richer label distribution than direct numerical regression, leading to more effective use of limited labels and better noise robustness.
@@
where $p_{i}$ is the predicted probability for the bucket $\mathbf{b}_{i}$ after softmax. These outputs form a distribution, from which we compute the expectation to obtain the final score as follows:
@@
where $y_{i}$ and the $\hat{y}_{i}$ are the ground truth label and the predicted score of sample $x_{i}$ , respectively. $|{\cal X}|$ is the number of samples in $\cal X$}. In DRILL, both the teacher and student models utilize the MAE loss function to incorporate supervised information, as detailed below.
Distillation for Semi-supervised Regression
Since using DDE alone is insufficient for effective representation learning, we reformulate the knowledge distillation process to enable the student model to learn more generalized features for improved performance in SSR. To build the distillation framework for SSR, we first apply weak augmentation to the data before feeding it into the teacher model, while strong augmentation is applied to the input of the student model. The predicted scores from the teacher model are used to guide the student by assigning corresponding pseudo-labels to the unlabeled nodes. As a result, the student model is trained on the entire dataset, including both labeled samples and unlabeled samples with pseudo-labels. The loss functions for the teacher and student models are thus defined separately as follows.
@@
where $\hat{y}_{i}^{\mathcal{T}}$ and $\hat{y}_{i}^{\mathcal{S}}$ are the predicted scores of the teacher model and the student model, respectively. To enable the student model to learn more accurate information for SSR, we define the loss $\mathcal{L}_{\text{DRILL}}$ as follows, based on the joint training of the teacher and student models.
$\mathcal{L}_{DRILL}=\mathcal{L}_{\mathcal{T}}+\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{Aligment}$ (7)
@@
The joint training mechanism enables the teacher model to generate high-quality pseudo labels from a small amount of labeled data, continuously improving the utilization of unlabeled data. The alignment loss $\mathcal{L}_{Alignment}$ minimizes the distance between the representations of the student and teacher models, typically implemented using Kullback–Leibler (KL) divergence or Mean Squared Error Loss (MSE).
Decoupling Distribution Alignment (DDA)

### v2

# Toward Robust Semi-supervised Regression via Dual-stream Knowledge Distillation

@@ Abstract
Semi-supervised regression (SSR), which aims to predict continuous scores of samples while reducing reliance on a large amount of labeled data, has recently received considerable attention across various applications, including computer vision, natural language processing, and audio and medical analysis. Existing SSR methods typically train models on scarce labeled data by introducing constraint-based regularization or ordinal ranking to reduce overfitting. However, these approaches fail to fully exploit the abundance of unlabeled samples. While consistency-driven pseudo-labeling methods attempt to incorporate unlabeled data, they are highly sensitive to pseudo-label quality and noisy predictions. To address these challenges, we introduce a Dual-stream Knowledge Distillation framework (DKD), which is specially designed for the SSR task to distill knowledge from both continuous-valued knowledge and distribution information, better preserving regression magnitude information and improving sample efficiency. Specifically, in DKD, the teacher is optimized solely with ground-truth labels for label distribution estimation, while the student learns from a mixture of real labels and teacher-generated pseudo targets on unlabeled data. The distillation design ensures the effective supervision transfer, allowing the student to leverage pseudo labels more robustly. Then, we introduce an advanced Decoupled Distribution Alignment (DDA) to align the target class and non-target class between teacher and student on the distribution, enhancing the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor. Extensive experiments were conducted on diverse datasets, including audio, text, image, and medical data. We show that DKD exhibits strong generalization capabilities and outperforms the best competing methods by 4.97% and 11.81% in terms of overall MAE and $R^{2}$, respectively.
1 Introduction
@@
Semi-supervised regression (SSR) aims to rank or predict continuous scores of samples while reducing reliance on a large amount of labeled data. It has been widely applied in various scenarios, including age estimation, clinical score prediction, and audio-based assessments, and other real-world applications where a large amount of labeled data across diverse are difficult to obtain Ren et al. (2022); Gui et al. (2024); Bao et al. (2024). With the rapid advancement of deep learning in recent years, regression tasks using deep learning models across various domains have attracted significant attention Yang et al. (2024); Van Engelen and Hoos (2020), especially in the SSR task. Some approaches directly employ deep learning models such as CNNs for images and RNNs for time series, optimizing them with loss functions like Mean Squared Error (MSE) or Mean Absolute Error (MAE) on the available labeled data Mohammadi Foumani et al. (2024); Yang et al. (2022); Zhang et al. (2017a). However, such a straightforward application struggles to learn effective patterns in scenarios with scarce labeled data, owing to the heavy reliance of deep learning models on large-scale annotations Pan et al. (2024); Yin et al. (2022). Therefore, direct regression (DR) based models typically yield suboptimal performance.
@@
To exploit the abundance of unlabeled data, several consistency-driven pseudo-labeling methods that assign the pseudo-labels on the unlabeled samples have been proposed Zhang et al. (2025); Zheng et al. (2025). Pseudo-labels are typically generated by training on the available labeled samples, and consistency regularization is often used to assist the optimization process and pseudo-labeling Zhao et al. (2025); Zhao and Wen (2025). While pseudo-labeling helps mitigate the challenge of limited labeled data, as some predicted incorrect labels are directly treated as true labels during training, and the model is likely to overfit to the wrong annotations. Therefore, the effectiveness of these approaches heavily relies on the quality of the pseudo-labels Jo et al. (2024); Yin et al. (2022), and assessing the reliability of the generated pseudo-labels remains a challenging task.
Figure 1: MAE and test SRCC curves for direct regression (DR), Rankup, UCVME, and DKD on the BVCC dataset. Rankup yields the least competitive results, as its pairwise ranking method may not fully capture the continuous relationships. Although UCVME improves performance via consistency-constrained pseudo-labels, it still falls short of DKD.
To address these challenges, we introduce an advanced decoupled distillation framework (DKD), specifically designed for SSR, to effectively exploit unlabeled samples. DKD distills the knowledge from both continuous-valued knowledge and distribution information in an end-to-end manner, better preserving regression magnitude information and improving sample efficiency. Specifically, DKD is implemented by a teacher model optimized solely with ground-truth labels for label distribution learning, producing informative predictions, while the student learns from a mixture of real labels and teacher-generated pseudo scores on unlabeled data, ensuring effective supervision transfer. The distillation allows the student to leverage pseudo labels more robustly, mitigating the interference from noisy pseudo-supervision in SSR. We further introduce an advanced decoupling distribution alignment (DDA) approach to distill the distribution information from the labeled trained teacher model, where we decouple the distribution into two parts for alignment, rather than aligning the distribution directly. DDA aligns the target and non-target distributions between the teacher and student in a decoupled manner, with an adaptive learnable weighting scheme, thereby enabling the student to better exploit unlabeled samples while reducing overfitting to noisy pseudo-labels and learning a more well-calibrated regression model. As shown in the Fig 1, the DKD can learn more robust and more generalizable knowledge representation provided by the mixture of labeled and pseudo-labeled samples provided by the teacher model. It outperforms the conventional method, direct regression (DR), and strong competing methods, including RankUp Huang et al. (2024), UCVME Dai et al. (2023a), achieving the best performance on both MAE and SRCC. In summary, this work makes the following main contributions.
  • We propose DKD, the first end-to-end decoupled distillation framework for semi-supervised regression, which effectively directly distills both continuous-valued knowledge and distribution information in an end-to-end manner, better preserving regression magnitude information and improving sample efficiency under limited labels.
@@
  • We further introduce a novel alignment method, block-aware decouple distribution alignment (DDA), which enables the student model to identify important classes, enhancing the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor.
@@
  • Extensive experiments were conducted on the datasets from various domains, including audio, text, image, and medical data. The experimental results demonstrate that DKD outperforms the state-of-the-art semi-supervised regression methods.
2 Related Works
2.1 Semi-supervised Regression
@@
The SSR methods can be roughly categorized into pseudo-label-based methods, consistency regularization, and contrastive-learning-based methods. Jo et al. Jo et al. (2024) propose an SSR framework that uses uncertainty estimation to guide pseudo-labeling, alongside a pseudo-label calibration approach that enhances pseudo-label quality by propagating information from labeled to unlabeled samples. Semi-supervised deep kernel learning Jean et al. (2018) minimizes predictive variance on unlabeled data through consistency regularization, leveraging the robust uncertainty quantification provided by Gaussian processes. Some contrastive-learning-based methods exploit the ranking information among samples as auxiliary for regression Qiao et al. (2022); Huang et al. (2024). For example, RankUp Huang et al. (2024) constructs pairwise ranking labels from continuous targets with an auxiliary ranking classifier. CLSS Dai et al. (2023a) and GCLSS Wang et al. (2025) facilitate contrastive regression by exploiting the ordinal relationships between unlabeled samples with a spectral seriation algorithm.
Although these methods have achieved remarkable success in SSR, most of them do not fully leverage the potential of abundant unlabeled data. Pseudo-labeling approaches incorporate unlabeled samples into training, but they are often sensitive to noisy pseudo-targets and may even degrade performance. In DKD, we propose an advanced decoupled representation distillation strategy that transfers knowledge from both continuous-value supervision and distributional information, enabling effective utilization of both pseudo-labeled and ground-truth labeled samples.
2.2 Distribution Alignment in Distillation
@@
Distribution alignment is a strategy used during student model training to match not only the outputs of the teacher model but also its underlying distributions. Existing approaches can be roughly categorized into logit-level alignment and feature-level alignment. Logit-level alignment aims to align the output distribution (e.g., softmax probabilities) of the teacher and student Hinton et al. (2015); Zhao et al. (2022); Zeng et al. (2025). For example, the multi-level logit distillation method Jin et al. (2023) employs multi-level prediction alignment to facilitate instance prediction learning in the student model. Curriculum temperature Li et al. (2023b) introduces temperature scheduling for dynamically adjusting softening intensity across training phases, which smooths gradient flow. The feature-level distribution alignment matches the representation, which is often done by maximum mean discrepancy, contrastive, and correlation-based loss Gou et al. (2021); Chen et al. (2021). Notable advancements in feature-level distillation encompass DiffKD Huang et al. (2023), training diffusion denoisers with teacher features to enforce attention map congruence in critical regions; WCoRD Chen et al. (2021), constructing global-local hybrid contrastive objectives with Wasserstein dual loss to align deep geometric structures. These alignment approaches have demonstrated effectiveness in knowledge transfer. Although there are some attempts by applying the decoupled distillation on the classification task Zhao et al. (2022); Li et al. (2023a); Feng et al. (2024). There is no work on semi-supervised regression that more emphasizes how pseudo-labels are leveraged effectively. Another difference of our method is that we incorporate the adaptive learnable weighting scheme specifically for the non-target alignment to further prevent the student from overfitting noisy pseudo-supervision. Meanwhile, the continuous-valued knowledge is incorporated into the distribution alignment on the basis of label distribution learning and alignment, to keep the regression magnitude information.
Figure 2: The Overview of DKD. (1) The input of DKD includes both labeled and unlabeled data. The teacher model is exclusively trained on the labeled data and is then used to generate pseudo labels for the unlabeled samples. The student model is trained on the entire dataset with a mixture of real labels and teacher-generated pseudo targets on unlabeled data. (2) Both the teacher and student are trained under a label-distribution formulation of regression, while continuous-valued distillation enforces consistency by minimizing the gap between their expected scores. (3) In the distribution distillation, DDA first identifies the target class to construct a binary probability distribution, then aligns the predicted distributions over classes between the teacher and student models separately for the target and non-target parts.
3 Methodology
@@
Notations. We assume the data set $\mathcal{X}$ consists of $n$ labeled samples $\mathcal{X}_{l}=\{{({x}_{i},y_{i})}_{i=1}^{n}\}$ and $m$ unlabeled samples $\mathcal{X}_{u}=\{{({x}_{i})}_{i=m}^{n+m}\}$, with instance $x\in\mathbb{R}$ for image, text, audio or tabular medical data. $\mathcal{Y}=\{y_{i}\}$ denotes the set of ground truth labels, where each label $y_{i}$ is a continuous value in $\mathbb{R}$.
@@
Problem Statement. Deep semi-supervised regression can be reformulated as given the training set $\mathcal{X}_{T}$ and $\mathcal{X}_{T}$ = $\mathcal{X}_{l}$, we aim to train a mapping function $f:\mathcal{X}_{u}\mapsto\mathcal{Y}$ which maps each input ${x}_{i}\in\mathcal{X}_{u}$ to a continuous scalar $y_{i}\in\mathbb{R}$. Semi-supervised learning explicitly incorporates unlabeled samples to complement limited labeled data during training. Therefore, in this case, the $\mathcal{X}_{T}$ consists of $\mathcal{X}_{l}$ and $\mathcal{X}_{u}$.
3.1 The Overview of DKD
@@
As shown in Fig. 2, the proposed DKD predicts scores by establishing a decoupled representation distillation framework, where the ground truth is transformed into a label distribution, effectively converting the regression task into a discrete distribution estimation problem for both teacher and student models. We then apply DKD over the label distribution and DDA for the distribution alignment to enable the student model to learn fine-grained and more generalized knowledge.
@@
This distillation process is particularly effective for the SSR task due to two key factors: (1) we reformulate the regression task as a distillation on both continuous-valued knowledge and distribution information in an end-to-end manner, which effectively mitigates the risk of the noise when integrating the pseudo labels, especially in the scenario that labeled samples are limited; and (2) we introduce a novel decoupled distribution alignment strategy for representation distillation over the classes, enabling the model to capture fine-grained features more effectively than conventional alignment methods. Finally, the model is jointly trained by combining the loss functions of the student and teacher, augmented with the proposed decoupled distribution alignment.
3.2 Distillation for Semi-supervised Regression
@@
Label Distribution Learning Conventional SSR methods typically predict the score by directly mapping the input sample ${x}_{i}$ to the target value $y$. However, this approach is highly sensitive to noise and irrelevant features, often resulting in poor generalization and a high risk of overfitting, especially when few labels are available. Therefore, following Gao et al. (2017); Xu et al. (2019), we define several classes $\mathcal{B}=\{{\mathbf{b}_{i}}\}_{i=1}^{L}$ based on the range of ground truth $y$, converting the regression task into a discrete distribution estimation problem. $\mathcal{B}$ can be intuitively viewed as a set of evenly spaced score bins (anchors) that discretize the continuous regression label range into $L$ classes, each covering an interval of width $c$, $\mathbf{b}_{i}=(i-1)\times c$, where each score is assigned to one corresponding class. Then, $y$ can be obtained by calculating the expectation as follows:
@@
where $\tilde{p}_{i}$ is the output probability of the model with the corresponding class $\mathbf{b}_{i}$}. Therefore, the mapping function $f$ in SSR is modified to learn a distribution $\mathcal{P}=\{p_{i}\}_{i=1}^{L}$ over predefined classes. By transforming the numerical regression into an interval prediction problem, it offers a smoother learning objective and richer label distribution than direct numerical regression Xu et al. (2019), leading to more effective use of limited labels and better noise robustness.
@@
where $p_{i}$ is the predicted probability for the class $\mathbf{b}_{i}$ after softmax. These outputs form a distribution, from which we compute the expectation to obtain the final score as follows:
@@
where $y_{i}$ and the $\hat{y}_{i}$ are the ground truth label and the predicted score of sample $x_{i}$ , respectively. $|{\cal X}|$ is the number of samples in $\cal X$}. In DKD, both the teacher and student models utilize the MAE loss function to incorporate supervised information, as detailed below.
Continuous-valued Knowledge Distillation Since using label distribution learning alone is insufficient for effective representation learning, we reformulate the knowledge distillation process to enable the student model to learn more generalized features for improved performance in SSR. To build the distillation framework for SSR, we first apply weak augmentation like random crop, flip, to the data before feeding it into the teacher model, while strong augmentation, like RandAugment Cubuk et al. (2020) is applied to the input of the student model. The detailed description of strong/weak augmentation on different domain data can be found in App. C.
The predicted scores from the teacher model are used to guide the student by assigning corresponding pseudo-labels to the unlabeled nodes. As a result, the student model is trained on the entire dataset, including both labeled samples and unlabeled samples with pseudo-labels. The loss functions for the teacher and student models are thus defined separately as follows.
@@
where $\hat{y}_{i}^{\mathcal{T}}$ and $\hat{y}_{i}^{\mathcal{S}}$ are the predicted scores of the teacher model and the student model, respectively. To enable the student model to learn more accurate information for SSR, we define the loss $\mathcal{L}_{\text{DKD}}$ as follows, based on the joint training of the teacher and student models.
$\mathcal{L}_{DKD}=\mathcal{L}_{\mathcal{T}}+\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{Aligment}$ (7)
@@
The joint training mechanism enables the teacher model to generate high-quality pseudo labels from a small amount of labeled data, continuously improving the utilization of unlabeled data. The alignment loss $\mathcal{L}_{Alignment}$ minimizes the distance between the representations of the student and teacher models, typically implemented using Kullback–Leibler (KL) divergence or MSE. In DKD, we align the label distributions and further introduce a decoupled distribution alignment strategy, which will be described in the following sections.
3.3 Decoupling Distribution Alignment (DDA)